what's the first five term of the sequence : tn = 3n-4
Answer by JBarnum(2146)

if n=1,2,3,4,5 (plug in the value for "n" and solve for "t")
then t=-1,1,5/3,2,11/5
if n=1:
t(1) = 3(1)-4
t=3-4
t=-1


if n=2:
t(2) = 3(2)-4
2t=6-4
2t=2
t=1